### v1

# MolTailor: Tailoring Chemical Molecular Representation to Specific Tasks via Text Prompts

Abstract
Deep learning is now widely used in drug discovery, providing significant acceleration and cost reduction. As the most fundamental building block, molecular representation is essential for predicting molecular properties to enable various downstream applications. Most existing methods attempt to incorporate more information to learn better representations. However, not all features are equally important for a specific task. Ignoring this would potentially compromise the training efficiency and predictive accuracy. To address this issue, we propose a novel approach, which treats language models as an agent and molecular pretraining models as a knowledge base. The agent accentuates task-relevant features in the molecular representation by understanding the natural language description of the task, just as a tailor customizes clothes for clients. Thus, we call this approach MolTailor. Evaluations demonstrate MolTailor’s superior performance over baselines, validating the efficacy of enhancing relevance for molecular representation learning. This illustrates the potential of language model guided optimization to better exploit and unleash the capabilities of existing powerful molecular representation methods. Our codes and appendix are available at https://github.com/SCIR-HI/MolTailor.
Introduction
In recent years, AI technology has been widely applied to various stages of drug design, such as compound synthesis and screening, etc. (Mamoshina et al. 2016). This has greatly improved the efficiency and reduced the cost of drug development. Molecular representation learning serves as the cornerstone for AI techniques in drug design, empowering a wide range of downstream tasks such as molecular property prediction and drug-drug interaction judgment (Xia et al. 2023). Molecular representations are essentially vector embeddings for molecules, analogous to word embeddings in NLP. The idea of encoding molecules as mathematical vectors dates back to the 1940s (Wiener 1947).
The development of molecular representation learning can be roughly divided into four phases. In the first phase, molecular representations were constructed through expert knowledge, which can be categorized into two types: descriptors (Wiener 1947) and fingerprints (Rogers and Hahn 2010). Descriptors focus on the inherent chemical characteristics of molecules, while fingerprints contain structural information. In the second phase, inspired by deep learning, some works (Gilmer et al. 2017; Yang et al. 2019; Song et al. 2020) started to learn molecular representations from labeled data using supervised learning. However, the scarcity of labeled data, due to the expensive experimental cost and time required, limits further improvements to model performance and generalization.
Figure 1: Most existing molecular pretraining models (e.g. Grover, MolCLR) attempt to encode as much molecular information as possible (e.g. various functional groups and molecular weight) into a vector to obtain general molecular representation. However, for specific downstream tasks (e.g. Lipo, predicting lipophilicity of compounds), features are not equally important (e.g. Sulfonamide and Carboxylic acid groups significantly increase the hydrophilicity, being more critical than the remaining groups). By understanding task descriptions, MolTailor adjusts the weights of different features in the representation to obtain task-specific molecular representation.
The success of pretrained language models in NLP (Devlin et al. 2018) has demonstrated the potential of self-supervised learning. Building upon this, in the third phase, some works pretrained sequence-based models to learn molecular representation by mimicking successful NLP pretrained language models. (Chithrananda, Grand, and Ramsundar 2020; Kim et al. 2021; Ross et al. 2022). Concurrently, others pretrained graph-based models (Hu et al. 2019; You et al. 2020). More recently, in the fourth phase, researchers are no longer limited to pretraining solely on the molecular data itself. Instead, they are attempting to incorporate additional information, such as knowledge graphs (Fang et al. 2022) and textual information (Zeng et al. 2022; Edwards et al. 2022; Su et al. 2022; Chen et al. 2023). Notably, molecule-text multimodal learning has received increasing attention recently and achieved promising results.
As introduced above, most existing works merely strive to inject more information into the representations without utilizing task information as prior knowledge, which could compromise the training efficiency, as shown in Fig. 1. Inspired by this observation, we propose a new method called MolTailor, where a language model acts as a tailor, adapting general molecular representations (ready-made clothing) into task-specific ones (customized clothing) based on user needs.
To achieve this goal, we adopt a dual-tower model structure with one language pretraining module and one molecular pretraining module, joint together through a cross-attention module. Meanwhile, we constructed a new pretraining task: Molecule-Text Multi-Task Regression (MT-MTR). The dataset for this task consists of (molecule, task description, regression labels) triplets. Here, the molecule is represented as a SMILES string (Weininger 1988), the task description is a text prompt describing molecular properties that are most helpful for solving the task, and the regression labels are values of features mentioned in the task description. The model needs to predict the regression labels based on the SMILES and task description. This pretraining task teaches the model to enhance the weights of task-relevant features in the molecular representation according to the task description. Our contributions can be summarized as follows:
  • We propose MolTailor, the first approach to generate task-specific molecular representations via text prompts, which provides a new perspective on text-molecule multimodal learning: not only injecting the knowledge in texts into molecular representations but also utilizing the reasoning capacity of language models.
  • We construct MT-MTR, a new molecule-text multimodal pretraining task. This task teaches the model the capabilities of instruction following and adapting molecular representations.
  • We comprehensively evaluate across eight subsets of the MoleculeNet, thereby demonstrating the effectiveness of task-specific molecular representation learning.
Related Work
Molecular pretraining models
Models can be roughly categorized into the following three types: sequence-based, graph-based, and external knowledge (Xia et al. 2023). 1. Sequence-based: Representing molecules as SMILES strings or other sequences and using language models as backbones for pretraining. SMILES-BERT (Wang et al. 2019) and ChemBERTa (Chithrananda, Grand, and Ramsundar 2020) use Masked Language Modeling (MLM) for pretraining tasks, while CHEM-BERT (Kim et al. 2021), ChemBERTa-2 (Ahmad et al. 2022) additionally incorporate property prediction tasks. MM-Deacon (Guo et al. 2021) uses contrastive learning with SMILES and INPAC as parallel inputs. 2. Graph-based: Representing molecules as graphs and using graph neural networks for pretraining. GROVER (Rong et al. 2020) and Mole-BERT (Xia et al. 2022) use Mask Component Modeling (MCM) for pretraining, while GraphCL (You et al. 2020) and MolCLR (Wang et al. 2022) employ contrastive learning. Denoising (Zaidi et al. 2022) and GeoSSL (Liu, Guo, and Tang 2022) are trained with denoising methods inspired by denoising diffusion models. 3. External knowledge: GraphMVP (Liu et al. 2021) and 3D Infomax (Stärk et al. 2022) incorporate 3D structures as supplementary information. KCL (Fang et al. 2022) uses knowledge graphs to enhance molecular representations.
Image-text multimodal pretraining models.
Models can be divided into dual-tower and single-tower structures based on architectures: 1. Dual-tower structure: ViLBERT (Lu et al. 2019) introduces the co-attentional transformer (Co-TRM) layer, which effectively integrates information from both modalities. CLIP (Radford et al. 2021) adopts contrastive learning for pretraining and achieves significant performance gains. The CoCa (Yu et al. 2022) model uses dual towers of an image encoder and a text decoder. 2. Single-tower structure: VisualBERT (Li et al. 2019) takes texts and images as a unified input for learning. ViLT (Kim, Son, and Kim 2021) adopts a patch projection approach, greatly improving the processing speed. BeiT-3 (Wang et al. 2023) uses distinct Feedforward Neural Network (FFN) layers for different modalities while sharing attention modules.
Molecule-text multimodal pretraining models.
Similarly, models can be divided into dual-tower and single-tower structures based on architectures: 1. Single-tower structure: Using a language model as the backbone for further pretraining. KV-PLM (Zeng et al. 2022) and MolXPT (Liu et al. 2023) inject SMILES into literature text by locating molecule names to obtain mixed corpora and conduct MLM and language pretraining (LM) respectively. MolT5 (Edwards et al. 2022) does replace corrupted spans pretraining on molecular and text data, meanwhile, uses mutual translation between molecules and textual descriptions as downstream tasks. Text+Chem T5 (Christofidellis et al. 2023) and ChatMol (Zeng et al. 2023) use multi-task learning for pretraining. GIMLET (Zhao et al. 2023) takes graphs and texts as a unified input and uses instruction-based supervised property prediction for pretraining. 2. Dual-tower structure: MoMu (Su et al. 2022) and MoleculeSTM (Liu et al. 2022) do contrastive learning on molecule-description pairs. CLAMP (Seidl et al. 2023) does contrastive learning on molecule-bioassay pairs. Additionally, ChemCrow (Bran et al. 2023) enhances large language models (LLMs) with chemical tools to accomplish real-world chemical tasks.
Methodology
Figure 2: Overview of the MolTailor framework. a) The construction process of the MT-MTR dataset. We obtain representative molecules from DrugBank (Wishart et al. 2018) and ChEBI (Hastings et al. 2016), and then use RDKit to calculate 209 properties for each molecule. For each molecule, we randomly sample 5-10 properties from the property set, use the property names to generate virtual task descriptions via GPT-3.5, and use the property values as regression labels. b) Model architecture of MolTailor. MolTailor consists of a language pretraining model (T-Encoder) and a molecular pretraining model (M-Encoder). The T-Encoder is divided into a unimodal part (for understanding task descriptions) and a multimodal part (for adjusting molecular representations). c) Internal structure of the Multimodal T-Encoder. It modifies the original Transformer Encoder Block to perform self-attention and cross-attention operations simultaneously: mapping the general molecular representation to obtain $K_{m}$ and $V_{m}$ vectors which are then concatenated with textual vectors $K_{t}$ and $V_{t}$. d) Pretraining task of MolTailor. The model needs to predict properties mentioned in the task description based on the molecule and text prompt. e) Downstream tasks of MolTailor. For a specific downstream task, we first generate the task description in the same format as pretraining via GPT-4 analysis, then take the SMILES and task description as input to predict labels for the corresponding task.
Fig. 2 presents an overview of our proposed approach. In this section, we introduce MolTailor in three aspects: 1. Construction of the MT-MTR corpus. 2. Model architecture of MolTailor. 3. Pretraining of MolTailor and application of MolTailor to downstream tasks.
Construction of MT-MTR
The construction process of the MT-MTR (Molecule-Text Multi-Task Regression) dataset consists of 3 main steps as shown in Fig. 2a:
Step 1: Obtain representative molecules. We obtain molecules from DrugBank (Wishart et al. 2018) and ChEBI (Hastings et al. 2016), then use RDKit [^1] https://www.rdkit.org to deduplicate and canonicalize, resulting in 55,759 valid SMILES.
Step 2: Calculate molecular properties. For each molecule obtained in the previous step, we use RDKit to calculate 209 properties [^2] RDKit’s Descriptor module provides 209 descriptors, which we uniformly refer to as “properties” in this paper for clarity.. Among these properties, some are continuous values while others are discrete values. We unify them into regression tasks.
Step 3: Obtain task descriptions and regression labels. For each molecule, we randomly sample 5-10 properties from the property set obtained in the previous step. The sampled property names are filled into the prompt template in Tab. 1 (Prompt for Pretraining) to obtain text input fed into GPT for generating virtual task description. Since GPT-4 and GPT-3.5 achieve similar performance on this task, we use the more cost-effective GPT-3.5 for generation in our experiments. The values of the sampled properties are used as the regression labels.
The resulting MT-MTR corpus contains (molecule, task description, regression labels) triplets. It is also worth noting that MT-MTR does not ask the model to predict all properties of the molecules; it only needs to predict properties mentioned in the text prompt. This aims to teach the model to generate molecule representations tailored to the text for improved predictions.
Template
Prompt
for
Pre-
training
As a seasoned expert in the field of chemistry, your task is to analyse a chemical task. And you found following properties of chemical compounds can help solve this task. Please summarize your analysis. The length should be less than 300 tokens.
Properties:
{Sampled Properties}.
Task analysis results:
Prompt
for
Down-
stream
Example:
{Example of task description from MT-MTR}
Property Names:
{209 property names from RDKit}
Please analyze the Task Name. When discussing the properties related to this task, list any properties that you think may be helpful for solving the task. Simply provide the analysis results directly in less than 400 tokens, referring to the example for guidance.
Table 1: Prompt template for interacting with GPT to generate task descriptions
Model architecture of MolTailor
As illustrated in Fig. 2b&c, MolTailor consists of a language pretraining model (T-Encoder) and a molecular pretraining model (M-Encoder). The language model is divided into a Unimodal Text Encoder (UT-Encoder) and a Multimodal Text Encoder (MT-Encoder) by introducing a few parameters. We treat the M-Encoder as a knowledge base and the T-Encoder as an agent. The agent obtains the final desired representation by understanding natural language and adjusting the well-initialized molecular representation from the knowledge base. The UT-Encoder captures semantic information and the MT-Encoder then tailors molecular representations based on understanding task descriptions.
Since any molecular pretraining model can serve as the M-Encoder in our method, in this section we mainly describe the Transformer Encoder Block (TEB), the fundamental component composing mainstream language models, as well as the internal structure of the MT-Encoder. Additionally, in our experiments, we tested two types of M-Encoder: CHEM-BERT (Kim et al. 2021) and ChemBERTa-2 (Ahmad et al. 2022).
Transformer Encoder Block.
As the basic building block of the Transformer (Vaswani et al. 2017) encoder, it consists of four main components: Multi-Head Self-Attention (MHA), Feed Forward Network (FFN), and Residual Connection with Layer Normalization (LN). For MHA, given hidden states $\boldsymbol{x}\in\mathbb{R}^{(bs\times n\times d)}$ where $bs$ indicates batch size, $n$ indicates sequence length, and $d$ indicates embedding dimension, it is first mapped to vectors $\boldsymbol{Q}$, $\boldsymbol{K}$ and $\boldsymbol{V}$ by matrices $\boldsymbol{W}_{Q},\boldsymbol{W}_{K},\boldsymbol{W}_{V}\in\mathbb{R}^{(d\times d)}$, keeping dimensions unchanged. Then, with $h$ heads, $d$ is split into $h$ parts. After that, vectors are transposed and reshaped to get $\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V}\in\mathbb{R}^{(bs\times h\times n\times d/h)}$. Finally, attention is computed:
$\mathrm{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\text{softmax}\left(\frac{\boldsymbol{Q}\boldsymbol{V}{{}^{\boldsymbol{T}}}}{\sqrt{d_{k}}}\right)$ (1)
$\mathrm{MHA}(\boldsymbol{x})=\mathrm{Attention}(f_{Q}^{r}(\boldsymbol{x}),f_{K}^{r}(\boldsymbol{x}),f_{V}^{r}(\boldsymbol{x}))$ (2)
where $f_{*}^{r}$ means $f^{reshape}(\boldsymbol{x}\boldsymbol{W}_{*})$. The output $\boldsymbol{x}^{*}\in\mathbb{R}^{(bs\times n\times d)}$ of the MHA goes through the FFN, residual connection and LN to obtain the final output $\boldsymbol{x}\in\mathbb{R}^{(bs\times n\times d)}$.
$\mathrm{FFN}(\boldsymbol{x})=\mathrm{ReLU}(\boldsymbol{x}\boldsymbol{W}_{1}+\boldsymbol{b}_{1})\boldsymbol{W}_{2}+\boldsymbol{b}_{2}$ (3)
$\boldsymbol{x}=\mathrm{LN}(\boldsymbol{x}+\mathrm{FFN}(\mathrm{MHA}(\boldsymbol{x})))$ (4)
Multimodal T-Encoder.
Since the task descriptions contain biomedical terminology, we chose PubMedBERT (Gu et al. 2021) as the backbone for the T-Encoder. Additionally, we also present in the appendix the results of using BioLinkBERT (Yasunaga, Leskovec, and Liang 2022). PubMedBERT contains 12 TEBs. We designate the first 9 layers as the UT-Encoder, and the remaining 3 layers as the MT-Encoder.
To build the MT-Encoder, we draw on previous work (Chen et al. 2022) to modify the original $\mathrm{MHA}$ to $\mathrm{MHA}^{*}$ that can simultaneously perform Self-Attention and Cross-Attention operations, by introducing very few parameters. Specifically, the MT-Encoder takes $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{m}$ as inputs, where $\boldsymbol{x}_{t}\in\mathbb{R}^{(bs\times n_{t}\times d_{t})}$ denotes the hidden states from the UT-Encoder, and $\boldsymbol{x}_{m}\in\mathbb{R}^{(bs\times n_{m}\times d_{m})}$ denotes the general molecular representation (i.e., the last hidden states from the M-Encoder). $\mathrm{MHA}^{*}$ computes as follows:
$\boldsymbol{Q}^{*}$ $=f^{r}(\boldsymbol{x}_{t}\boldsymbol{W}_{Q}^{t})$ (5)
$\boldsymbol{K}^{*}$ $=f^{r}([\boldsymbol{x}_{t}\boldsymbol{W}_{K}^{t},\boldsymbol{x}_{m}\boldsymbol{W}_{K}^{m}])$
$\boldsymbol{V}^{*}$ $=f^{r}([\boldsymbol{x}_{t}\boldsymbol{W}_{V}^{t},\boldsymbol{x}_{m}\boldsymbol{W}_{V}^{m}])$
$\mathrm{MHA}^{*}(\boldsymbol{x}_{t},\boldsymbol{x}_{m})$ $=\mathrm{Attention}(\boldsymbol{Q}^{*},\boldsymbol{K}^{*},\boldsymbol{V}^{*})$
where $f^{r}$ denotes reshape function, [*]delimited-[][*][ * ] denotes concatenation along the dimension corresponding to the sequence length $n$. The output of the MT-Encoder Block is computed:
$\boldsymbol{x}^{*}=\mathrm{LN}(\boldsymbol{x}_{t}+\mathrm{FFN}(\mathrm{MHA}^{*}(\boldsymbol{x}_{t},\boldsymbol{x}_{m})))$ (6)
Chen et al. (2022) formally proves that $\mathrm{MHA}^{*}(\boldsymbol{x})$ is equivalent to a weighted average of self-attention and cross-attention:
$\mathrm{MHA}^{*}(\boldsymbol{x}_{t},\boldsymbol{x}_{m})$ $=(1-\lambda(\boldsymbol{x}_{t}))\mathrm{MHA}(\boldsymbol{x}_{t})$ (7)
$+\lambda(\boldsymbol{x}_{t})\mathrm{MHA}(\boldsymbol{Q}_{t},\boldsymbol{K}_{m},\boldsymbol{V}_{m})$
Pretraining and Downstream Application
Pretraining on MT-MTR.
We use MT-MTR as the pretraining objective . In detail, given SMILES ($\boldsymbol{z}_{m}$) and task description ($\boldsymbol{z}_{t}$) as input, the model predicts regression labels $\boldsymbol{y}\in\mathbb{R}^{(1\times 209)}$, as displayed in Fig. 2d. The pretraining loss uses MSE and can be formulated as:
$\mathcal{L}=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{1}{\mathrm{count}(\boldsymbol{m})}\sum_{i=1}^{M}m_{ij}(y_{ij}-\hat{y}_{ij})^{2}\right)$ (8)
where $N$ is the number of samples, $\boldsymbol{m}\in\mathbb{R}^{(1\times 209)}$ is a 0-1 vector, in which the value 1 indicates the existence of the corresponding regression label, while 0 indicates its absence, $\mathrm{count}(\boldsymbol{m})$ means the number of valid labels, $M$ is the number of all properties (here $M=209$), $y_{ij}$ is the ground truth regression label, and $\hat{y}_{ij}$ is the predicted value.
Downstream Application.
When applying MolTailor to downstream tasks, we use GPT-4 to analyze the specific task and then generate the corresponding task description, as shown in Fig. 2e. We use the template in Tab. 1 to prompt GPT to generate descriptions that mimic the format of those in the pretraining corpus. At the same time, we restrict it to select task-relevant features from the set of properties supported by RDKit. Then, we feed the generated analysis and SMILES into MolTailor to obtain the task-specific molecular representation:
$\boldsymbol{z}=\mathrm{MolTailor}\left(\boldsymbol{z}_{m}|\boldsymbol{z}_{t}\right)$ (9)
where $\boldsymbol{z}$ denotes the task-specific molecular representation. Finally, $\boldsymbol{z}$ goes through a prediction head to predict the labels corresponding to the task.
Experiments
In this section, we conduct comprehensive experiments to demonstrate the efficacy of our proposed method. The experiments are designed to analyze the method by addressing the following six key questions:
Q1: Are the task-specific molecular representations generated by MolTailor better than general representations? Q2: Can MolTailor achieve performance improvements on different M-Encoders? Q3: How do the task descriptions influence the effectiveness of pretraining? Q4: How do different text prompts affect MolTailor? Q5: When MolTailor achieves better performance, are the task-relevant features in the molecular representations enhanced? Q6: Does MolTailor pay attention to the key information in both the molecules and text prompts?
Experimental Setup
Pretraining Corpus.
We use MT-MTR corpus for pretraining, which contains 55,759 triples. Moreover, we present the data overlap analysis results in the appendix.
Downstream Datasets.
We select 8 representative tasks from MoleculeNet (Wu et al. 2018) for experiments, which consist of 4 classification tasks (BBBP, ClinTox, HIV, Tox21) and 4 regression tasks (ESOL, FreeSolv, Lipophilicity, QM8), covering physiology, biophysics, physical chemistry, and quantum mechanics. Following Wu et al. (2018), each task uses the recommended splitting method to divide data into training/validation/test sets with a ratio of 8:1:1.
Baselines.
We adopt the following four types of baselines:
  • Molecular Fingerprints: MACCSFP (Durant et al. 2002) encodes molecules based on substructures, RDKitFP (O’Boyle and Sayle 2016) encodes molecules based on topology or path, and MorganFP (Rogers and Hahn 2010) encodes molecular environment and structure starting from atoms within a radius.
  • Sequence-based Methods: BioLinkBERT(Yasunaga, Leskovec, and Liang 2022), PubMedBERT (Gu et al. 2021); ChemBERTa-2 (Ahmad et al. 2022), and CHEM-BERT (Kim et al. 2021).
  • Graph-based Methods: Grover (Rong et al. 2020), MolCLR (Wang et al. 2022), Mole-BERT (Xia et al. 2022), and Uni-Mol (Zhou et al. 2023).
  • Multimodal Methods: KCL (Fang et al. 2022), KV-PLM (Zeng et al. 2022), MolT5 (Edwards et al. 2022), and MoMu (Su et al. 2022).
We additionally construct Random and RDKit-DP as baselines for comparison, where Random refers to random vectors, and RDKit-DP consists of the 209 molecular properties calculated by RDKit.
Evaluation Methodology.
To better evaluate the molecular representations learned by different models, we conduct experiments in linear probe setting. As a result, the baseline experimental results reported in this paper may differ from the original papers. We freeze the model parameters to extract representations for downstream tasks. These extracted representations are then mapped to labels through a learnable linear layer.
Following the recommendation of Wu et al. (2018), we use ROC-AUC as the evaluation metric for classification tasks. For the regression task qm8, we use MAE, and for other regression tasks, we use RMSE. To ensure fairness, we use Optuna (Akiba et al. 2019) to search 10 learning rates (LRs) for each model. We repeat each task 3 times and report the mean and standard deviation. Due to space limitations, the standard deviations are included in the appendix.
Implementation Details.
For the pretraining phase, we employ the AdamW optimizer, complemented by a linear learning rate scheduler. We set the LR at 5.5e-5 and use a warmup ratio of 0.1. The training is conducted 50 epochs with a batch size of 64, utilizing two A100-SXM4-80GB GPUs. For the downstream tasks, we opt for the Adam optimizer and leverage Optuna for hyperparameter tuning, conducting 10 trials to identify the optimal LR within the range of 1e-5 to 1e-2 for each model on every task. The optimal LR is determined based on the performance of the validation set. We train our model using a batch size of 64 on a single GeForce RTX 2080 Ti GPU, employing an early stop mechanism with a patience setting of 3 and limiting the training to a maximum of 50 epochs.
Models Classification (ROC-AUC) Regression (RMSE / MAE) Params
Dataset BBBP ClinTox HIV Tox21 ESOL FreeSolv Lip QM8
#Molecules 2039 1478 41127 7831 1128 642 4200 21786
#Split Scaffold Random Scaffold Random Random Random Random Random
#Tasks 1 2 1 12 1 1 2 16
Random 48.38 56.01 49.54 51.11 3.3358 5.4831 1.3813 0.0320 -
RDKit-DP 78.25 67.36 70.85 65.61 4.8940 2.8068 0.9963 0.0202 -
RDKit-FP 87.65 57.13 78.66 76.14 1.0830 2.0725 0.9007 0.0181 -
MACCS-FP 81.64 83.05 77.53 77.27 1.0833 1.9086 0.9886 0.0196 -
Morgan-FP 82.73 72.61 82.65 75.29 1.2413 2.1896 0.8196 $0$
Grover 79.83 87.75 77.47 79.61 0.8977 1.9041 0.8301 0.0184 107M
MolCLR 81.27 78.15 71.48 75.61 1.3421 3.0436 1.0448 0.0219 2M
Mole-BERT 82.70 81.82 79.35 84.20 1.1379 2.3626 0.8316 0.0221 2M
Uni-Mol 79.52 88.65 74.18 78.08 1.0509 2.6913 1.0363 0.0219 48M
BioLinkBERT 83.81 87.75 71.24 73.81 1.1739 3.1350 1.0589 0.0234 108M
PubMedBERT 89.10 84.29 72.30 73.77 1.0715 2.5999 1.0851 0.0232 108M
ChemBERTa-2 84.70 84.21 78.88 80.75 0.8103 1.8439 0.7948 0.0191 3M
CHEM-BERT 84.10 93.80 76.99 80.54 0.7973 2.0214 0.8571 0.0215 51M
KCL 76.86 60.80 68.48 74.98 0.8728 2.7615 0.9868 0.0225 1M
KV-PLM 86.36 81.20 73.52 74.62 1.1785 2.8840 1.1004 0.0233 109M
MoMu-ME 80.41 67.99 71.91 74.75 1.4135 2.3229 0.9835 0.0222 2M
MoMu-TE 82.24 81.94 67.88 73.07 1.2562 3.1480 1.0885 0.0250 109M
$*{}^{*}$ 84.65 85.95 76.42 80.32 0.7128 1.7826 0.7848 0.0185 112M
MolTailor 81.15 92.37 77.42 80.67 0.7234 1.7881 0.8107 0.0196 161M
Table 2: Evaluation results of MolTailor and baselines under the linear probe setting on MoleculeNet. Here, ME stands for Molecule Encoder and TE stands for Text Encoder in MoMu-ME/TE. $*{}^{*}$ denotes using ChemBERTa-2 as the M-Encoder, while MolTailor indicates using CHEM-BERT. The table shows the average performance over 3 runs. Standard deviations are omitted due to space limitations but can be viewed in the appendix. Results in bold indicate the best performance. Green highlights cases where MolTailor outperforms the backbone, while yellow indicates the opposite.
Performance Comparison (Q1 & Q2)
Q1: We evaluate whether molecular representations enhanced by task descriptions could improve performance across 8 tasks. Tab. 2 shows that MolTailor achieves performance gains over the backbone model on the 4 regression tasks, and notably attains state-of-the-art (SOTA) results on ESOL, FreeSolv, and Lipophilicity datasets. On the 4 classification tasks, MolTailor’s performance is inconsistent, with gains on HIV and Tox21 but losses on the other 2 datasets when using CHEM-BERT as backbone. We hypothesize this may be because the pretraining task favors regression, which could inhibit performance on classification. And this implies classification and regression tasks may require different optimizing approaches. Additionally, we test converting MT-MTR to MT-MTC for pretraining but see minimal improvement, which indicates this is not due to the form of the task.
Q2: We experiment on two backbones, ChemBERTa-2 and CHEM-BERT. MolTailor achieves similar gains on top of both backbones. Notably, the performance differences between the backbones are also mirrored in the respective MolTailor variants. This demonstrates MolTailor’s transferability and ability to inherit strengths of different backbones. The results support that further gains may be achievable by transferring MolTailor to new state-of-the-art single-modality models.
Alation Study (Q3 & Q4 & Q5)
Q3. To evaluate the influence of task descriptions on pretraining in MT-MTR, we first remove the task descriptions to obtain the dataset $*{}^{*}$, then pretrain CHEM-BERT and PubMedBERT on $*{}^{*}$ to get $*{}^{*}$ and $*{}^{*}$ models. Next, we concatenate the task descriptions after the SMILES strings to construct dataset ${}^{\dagger}$, and pretrain PubMedBERT on ${}^{\dagger}$ to obtain ${}^{\dagger}$. As shown in Tab. 3, $*{}^{*}$ improves over untrained CHEM-BERT on regression tasks but drops on classification, indicating MTR as a pretraining task benefits downstream regression but negatively impacts classification. Also, ${}^{\dagger}$ over $*{}^{*}$ and MolTailor over CHEM-BERT both demonstrate further regression performance gains but classification performance drops. This shows introducing task descriptions helps models better learn from the data. That is, if constructing new labels can improve model performance on classification tasks, then supplementing task descriptions can further amplify such gains.
Models Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
CHEM-BERT 83.86 0.9243
$*{}^{*}$ 83.11 0.9209
PubMedBERT 79.87 1.1949
$*{}^{*}$ 81.67 0.9148
${}^{\dagger}$ 81.08 0.9131
MolTailor 82.09 0.8475
Table 3: Experimental results for Q3. For these experiments, the overlapping data are excluded, and all models are pretrained for only 20 epochs.
Q4. We replace the task-specific descriptions generated via GPT-4 analysis with the irrelevant text “to be or not to be, this is the question.” as noise prompts. Results in Tab. 4 show degraded performance with noise prompts, indicating information in the task descriptions does help the model generate better molecular representations. However, it should be noted that the performance drop is slight, suggesting textual information is not as important as expected in the process of generating representations. Further investigation into this phenomenon is needed.
Prompt Types Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
Origin 82.92 0.8348
Noise 82.77 0.8541
Table 4: Experimental results for Q4.
Q5. Earlier experiments have shown that task-specific molecular representations outperform general ones. Now, we aim to test if task-relevant features are more pronounced in these task-specific representations. Our theory is simple: if a feature is more prominent in the representation, it should lead to better predictions for that feature compared to a less distinct representation.
For our experiment, we chose the ESOL dataset and four models: KV-PLM, CHEM-BERT, MolTailor, and MolTailor (noise), where the latter refers to a version using a noise prompt. We calculated 209 properties for each molecule in ESOL, creating a new dataset named ESOL-MTR, comprising pairs of SMILES strings and their 209 regression labels. We trained all four models on ESOL-MTR and evaluated their ability to predict the 209 properties of the ESOL molecules. From these, we focused on 4 properties for Fig. 3. We calculated the Normalized RMSE by subtracting the average RMSE of the four models on a task from the original RMSE and then dividing it by the standard deviation. The four properties we selected include two mentioned in the task description (MolWt and FractionCSP3), one relevant but unmentioned (EState_VSA3), and one unrelated and unmentioned (Kappa1). Our analysis led to two main observations:
  • For the properties mentioned in the task description, MolTailor outperforms MolTailor (noise), confirming that these mentioned properties are indeed emphasized.
  • We find prompts influence both mentioned and unmentioned properties. “EState_VSA3” is a relevant but unmentioned property shows performance gains. In contrast, “Kappa1” is an irrelevant and unmentioned property performance decline. The results show that by understanding text prompts, the model can attend to properties not explicitly stated.
Figure 3: Normalized RMSE of four models on ESOL task and associated/unassociated proterties. The x-axis shows the names of tasks including ESOL and four properties, of which the first three are related to ESOL, while the last is not. The y-axis plots N-RMSE for the four models on each task, with lower values indicating better performance.
Case Study (Q6)
Q6. We analyze the attention matrices from the last layers of the UT-Encoder and MT-Encoder to investigate whether MolTailor pays attention to key information. The analysis results are shown in Fig. 4. Specifically, we select ESOL as a representative task related to solubility. If the model attends to information like molecular weight and polar functional groups critical for determining solubility, it suggests key information is captured.
After training on ESOL, we input SMILES from the test set and task description into the model to obtain attention matrices $\boldsymbol{M}_{UT}$ and $\boldsymbol{M}_{MT}$. Since we use the vector corresponding to the “[CLS]” token for prediction, we believe the larger the attention value between “[CLS]” and a token in $\boldsymbol{M}_{UT}$ or $\boldsymbol{M}_{MT}$, the more the model focuses on that information.
In Fig. 4, tokens highlighted in red in the top 5 most attended by the UT/MT-Encoders relate to molecular weight or polar functional groups, indicating the model does pay attention to key textual information. Also, the two molecular graphs plotted using the most attended molecular tokens by the MT-Encoder under different prompts show the original prompt leads to more attention on polar functional groups like ester and ketone, meaning critical information is better captured.
Figure 4: Visualization of the attention, answering Q6. The two molecular graphs on the left show MolTailor’s attention over the molecules under different prompts. The text on the right shows which input tokens from the original prompt MolTailor pays most attention to.
Conclusion and Future Work
Overall, this work provides a new perspective on molecular representation learning, not only striving to include more comprehensive information in the representations, but also combining contextual information and obtaining molecular representations more suitable for specific tasks through tailoring. In this work, we not only utilize the knowledge implied in the text modality, more importantly, we try to leverage the excellent reasoning ability of language models, which will become more important in the age of large models.
In the future, we first plan to explore new pretraining tasks that can stably improve model performance on both classification and regression tasks. Secondly, we will explore how to build molecular-text multimodal models based on large language models. Finally, we look forward to in-depth collaborations with domain experts to apply the molecular-text multimodal methods to actual production problems.
Acknowledgements
We thank the reviewers for their detailed and constructive suggestions and for the affirmation of our paper. Meanwhile, we gratefully acknowledge the support of the National Key R&D Program of China (2021ZD0113302), the National Natural Science Foundation of China Youth Fund (62206079), and the Heilongjiang Provincial Natural Science Foundation of China (YQ2022F006). We also appreciate the support from Du Xiaoman (Beijing) Science Technology Co., Ltd. @ on our research.

### v2

@@
 This illustrates the potential of language model guided optimization to better exploit and unleash the capabilities of existing powerful molecular representation methods.
-Our codes and appendix are available at https://github.com/SCIR-HI/MolTailor.
+Our code is available at https://github.com/SCIR-HI/MolTailor.
 Introduction
@@
 The dataset for this task consists of (molecule, task description, regression labels) triplets.
-Here, the molecule is represented as a SMILES string (Weininger 1988), the task description is a text prompt describing molecular properties that are most helpful for solving the task, and the regression labels are values of features mentioned in the task description.
+Here, the molecule is represented as a SMILES string (Weininger 1988), the task description is a text prompt describing molecular properties that are most helpful for solving the task, and the regression labels are values of properties mentioned in the task description.
 The model needs to predict the regression labels based on the SMILES and task description.
-This pretraining task teaches the model to enhance the weights of task-relevant features in the molecular representation according to the task description.
+This pretraining task teaches the model to enhance the weights of task-relevant properties in the molecular representation according to the task description.
 Our contributions can be summarized as follows:
@@
 Related Work
-Molecular pretraining models
+Molecular Pretraining Models.
 Models can be roughly categorized into the following three types: sequence-based, graph-based, and external knowledge (Xia et al.
@@
 Step 2: Calculate molecular properties.
-For each molecule obtained in the previous step, we use RDKit to calculate 209 properties [^2] RDKit’s Descriptor module provides 209 descriptors, which we uniformly refer to as “properties” in this paper for clarity..
+For each molecule obtained in the previous step, we use RDKit to calculate 209 properties [^2] RDKit’s “Descriptor” module provides 209 descriptors, which we uniformly refer to as “properties” in this paper for clarity..
 Among these properties, some are continuous values while others are discrete values.
@@
 Simply provide the analysis results directly in less than 400 tokens, referring to the example for guidance.
-Table 1: Prompt template for interacting with GPT to generate task descriptions
+Table 1: Prompt templates for interacting with GPT to generate task descriptions.
 Model Architecture of MolTailor
@@
 $\mathrm{MHA}^{*}(\boldsymbol{x}_{t},\boldsymbol{x}_{m})$ $=\mathrm{Attention}(\boldsymbol{Q}^{*},\boldsymbol{K}^{*},\boldsymbol{V}^{*})$
-where $f^{r}$ denotes reshape function, [*]delimited-[][*][ * ] denotes concatenation along the dimension corresponding to the sequence length $n$. The output of the MT-Encoder Block is computed:
+where $f^{r}$ denotes reshape function, []delimited-[][*][ ∗ ] denotes concatenation along the dimension corresponding to the sequence length $n$. The output of the MT-Encoder Block is computed:
 $\boldsymbol{x}^{*}=\mathrm{LN}(\boldsymbol{x}_{t}+\mathrm{FFN}(\mathrm{MHA}^{*}(\boldsymbol{x}_{t},\boldsymbol{x}_{m})))$ (6)
@@
 1 to prompt GPT to generate descriptions that mimic the format of those in the pretraining corpus.
-At the same time, we restrict it to select task-relevant features from the set of properties supported by RDKit.
+Notably, we restrict it such that it can only select the names of properties from the set supported by RDKit.
 Then, we feed the generated analysis and SMILES into MolTailor to obtain the task-specific molecular representation:
@@
 Q4: How do different text prompts affect MolTailor?
-Q5: When MolTailor achieves better performance, are the task-relevant features in the molecular representations enhanced?
+Q5: When MolTailor achieves better performance, are the task-relevant properties in the molecular representations enhanced?
 Q6: Does MolTailor pay attention to the key information in both the molecules and text prompts?
@@
 MoMu-TE 82.24 81.94 67.88 73.07 1.2562 3.1480 1.0885 0.0250 109M
-$*{}^{*}$ 84.65 85.95 76.42 80.32 0.7128 1.7826 0.7848 0.0185 112M
+MolTailor 84.65 85.95 76.42 80.32 0.7128 1.7826 0.7848 0.0185 112M
 MolTailor 81.15 92.37 77.42 80.67 0.7234 1.7881 0.8107 0.0196 161M
@@
 Here, ME stands for Molecule Encoder and TE stands for Text Encoder in MoMu-ME/TE.
-$*{}^{*}$ denotes using ChemBERTa-2 as the M-Encoder, while MolTailor indicates using CHEM-BERT.
+MolTailor denotes using ChemBERTa-2 as the M-Encoder, while MolTailor indicates using CHEM-BERT.
 The table shows the average performance over 3 runs.
 Standard deviations are omitted due to space limitations but can be viewed in the appendix.
-Results in bold indicate the best performance.
-Green highlights cases where MolTailor outperforms the backbone, while yellow indicates the opposite.
+Results in bold indicate state-of-the-art (SOTA) performance, while the underlined results show that the model outperforms the backbone used in the M-Encoder.
 Performance Comparison (Q1 & Q2)
@@
 On the 4 classification tasks, MolTailor’s performance is inconsistent, with gains on HIV and Tox21 but losses on the other 2 datasets when using CHEM-BERT as backbone.
-We hypothesize this may be because the pretraining task favors regression, which could inhibit performance on classification.
-And this implies classification and regression tasks may require different optimizing approaches.
-Additionally, we test converting MT-MTR to MT-MTC for pretraining but see minimal improvement, which indicates this is not due to the form of the task.
+We hypothesize that the difference in model performance on classification and regression tasks is due to the pretraining data benefiting the regression task more.
+Meanwhile, we experimented with converting the regression pretraining task into a classification form.
+The results were still similar, eliminating the influence of the form of the pretraining task on downstream tasks.
 Q2: We experiment on two backbones, ChemBERTa-2 and CHEM-BERT.
@@
 The results support that further gains may be achievable by transferring MolTailor to new state-of-the-art single-modality models.
-Alation Study (Q3 & Q4 & Q5)
+Ablation Study (Q3 & Q4 & Q5)
 Q3.
-To evaluate the influence of task descriptions on pretraining in MT-MTR, we first remove the task descriptions to obtain the dataset $*{}^{*}$, then pretrain CHEM-BERT and PubMedBERT on $*{}^{*}$ to get $*{}^{*}$ and $*{}^{*}$ models.
-Next, we concatenate the task descriptions after the SMILES strings to construct dataset ${}^{\dagger}$, and pretrain PubMedBERT on ${}^{\dagger}$ to obtain ${}^{\dagger}$.
+To evaluate the influence of task descriptions on pretraining in MT-MTR, we first remove the task descriptions to obtain the dataset MT-MTR, then pretrain CHEM-BERT and PubMedBERT on MT-MTR to get CHEM-BERT and PubMedBERT models.
+Next, we concatenate the task descriptions after the SMILES strings to construct dataset MT-MTR, and pretrain PubMedBERT on MT-MTR to obtain PubMedBERT.
 As shown in Tab.
-3, $*{}^{*}$ improves over untrained CHEM-BERT on regression tasks but drops on classification, indicating MTR as a pretraining task benefits downstream regression but negatively impacts classification.
-Also, ${}^{\dagger}$ over $*{}^{*}$ and MolTailor over CHEM-BERT both demonstrate further regression performance gains but classification performance drops.
+3, CHEM-BERT improves over untrained CHEM-BERT on regression tasks but drops on classification, indicating MTR as a pretraining task benefits downstream regression but negatively impacts classification.
+Also, PubMedBERT over PubMedBERT and MolTailor over CHEM-BERT both demonstrate further regression performance gains but classification performance drops.
 This shows introducing task descriptions helps models better learn from the data.
 That is, if constructing new labels can improve model performance on classification tasks, then supplementing task descriptions can further amplify such gains.
-Models Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
-CHEM-BERT 83.86 0.9243
-$*{}^{*}$ 83.11 0.9209
-PubMedBERT 79.87 1.1949
-$*{}^{*}$ 81.67 0.9148
-${}^{\dagger}$ 81.08 0.9131
-MolTailor 82.09 0.8475
-Table 3: Experimental results for Q3. For these experiments, the overlapping data are excluded, and all models are pretrained for only 20 epochs.
 Q4.
-We replace the task-specific descriptions generated via GPT-4 analysis with the irrelevant text “to be or not to be, this is the question.” as noise prompts.
+We replace the task-specific descriptions generated via GPT-4 with the irrelevant text “to be or not to be, this is the question.” as noise prompts.
 Results in Tab.
@@
 Further investigation into this phenomenon is needed.
-Prompt Types Classificationnormal-$\uparrow$ Regressionnormal-$\downarrow$
+[…]
+CHEM-BERT 83.86 0.9243
+CHEM-BERT 83.11 0.9209
+PubMedBERT 79.87 1.1949
+PubMedBERT 81.67 0.9148
+PubMedBERT 81.08 0.9131
+MolTailor 82.09 0.8475
+Table 3: Experimental results for Q3.
+CHEM-BERT and PubMedBERT use their original weights when evaluated on downstream tasks.
+The remaining methods are pretrained on MT-MTR and its variants, with overlapping molecules between pretraining and downstream data removed.
+Additionally, all methods are pretrained for only 20 epochs.
+[…]
 Origin 82.92 0.8348
@@
 Table 4: Experimental results for Q4.
+Figure 3: Performance of three methods on ESOL and molecular properties prediction tasks for Q5.
+The x-axis shows the task names, the y-axis shows the normalized RMSE, with lower values indicating better performance.
+Of the four molecular properties, the first three are related to the ESOL task, while the last one is opposite.
 Q5.
-Earlier experiments have shown that task-specific molecular representations outperform general ones.
-Now, we aim to test if task-relevant features are more pronounced in these task-specific representations.
-Our theory is simple: if a feature is more prominent in the representation, it should lead to better predictions for that feature compared to a less distinct representation.
-For our experiment, we chose the ESOL dataset and four models: KV-PLM, CHEM-BERT, MolTailor, and MolTailor (noise), where the latter refers to a version using a noise prompt.
-We calculated 209 properties for each molecule in ESOL, creating a new dataset named ESOL-MTR, comprising pairs of SMILES strings and their 209 regression labels.
-We trained all four models on ESOL-MTR and evaluated their ability to predict the 209 properties of the ESOL molecules.
-From these, we focused on 4 properties for Fig.
+The experiments in Tab.
+2 show that the molecule representations generated by MolTailor attained better performance.
+However, are the obtained representations more task-specific, meaning the task-relevant properties in the representations are enhanced?
+To validate this, we designed the following experiment: we used the representations generated by different methods to predict task-related and task-unrelated properties.
+If the expectations are met, then the representations enhanced by MolTailor should have better performance in predicting task-related molecular properties.
+We conduct experiments on the ESOL dataset.
+For the models, we use MolTailor and its corresponding backbone CHEM-BERT.
+We also test the performance of MolTailor using irrelevant text as the prompt, denoted as MolTailor(noise).
+In detail, we first use RDKit to compute the properties of the molecules in the ESOL dataset, obtaining the ESOL-MTR dataset.
+Then, we randomly split the dataset into 8:1:1 for training, validation, and testing.
+The results are shown in Fig.
 3.
-We calculated the Normalized RMSE by subtracting the average RMSE of the four models on a task from the original RMSE and then dividing it by the standard deviation.
-The four properties we selected include two mentioned in the task description (MolWt and FractionCSP3), one relevant but unmentioned (EState_VSA3), and one unrelated and unmentioned (Kappa1).
-Our analysis led to two main observations:
-  • For the properties mentioned in the task description, MolTailor outperforms MolTailor (noise), confirming that these mentioned properties are indeed emphasized.
-  • We find prompts influence both mentioned and unmentioned properties.
-“EState_VSA3” is a relevant but unmentioned property shows performance gains.
-In contrast, “Kappa1” is an irrelevant and unmentioned property performance decline.
-The results show that by understanding text prompts, the model can attend to properties not explicitly stated.
-Figure 3: Normalized RMSE of four models on ESOL task and associated/unassociated proterties.
-The x-axis shows the names of tasks including ESOL and four properties, of which the first three are related to ESOL, while the last is not.
-The y-axis plots N-RMSE for the four models on each task, with lower values indicating better performance.
+We present the performance of the three methods on the ESOL task and in predicting four molecular properties.
+Among the four selected properties, two have names mentioned in the prompt and are relevant to the ESOL task (MolWt and FractionCSP3), one is not mentioned but is relevant (EState_VSA3), and one is neither mentioned nor relevant (Kappa1).
+We use RMSE as the evaluation metric.
+For better visualization, we normalize the RMSE under each task using the formula: Normalized $\text{Normalized RMSE}=\tfrac{\text{RMSE}-\text{mean}(\text{RMSE})}{\text{std}(\text{RMSE})}$.
+The results show that: a) Compared to CHEM-BERT, MolTailor does focus more on the properties mentioned in the prompt, validating that the obtained representations are indeed task-specific.
+b) MolTailor not only pays more attention to the relevant properties explicitly stated in the prompt, but also to the unstated yet still relevant properties.
+Meanwhile, it decreases attention on the unrelated, unmentioned properties.
+This demonstrates the method’s generalization capability.
+c) The declined performance of MolTailor(noise) in property prediction indicates that the prompt does help the model attend to critical information.
 Case Study (Q6)
 Q6.
-We analyze the attention matrices from the last layers of the UT-Encoder and MT-Encoder to investigate whether MolTailor pays attention to key information.
-The analysis results are shown in Fig.
-4.
-Specifically, we select ESOL as a representative task related to solubility.
-If the model attends to information like molecular weight and polar functional groups critical for determining solubility, it suggests key information is captured.
-After training on ESOL, we input SMILES from the test set and task description into the model to obtain attention matrices $\boldsymbol{M}_{UT}$ and $\boldsymbol{M}_{MT}$.
-Since we use the vector corresponding to the “[CLS]” token for prediction, we believe the larger the attention value between “[CLS]” and a token in $\boldsymbol{M}_{UT}$ or $\boldsymbol{M}_{MT}$, the more the model focuses on that information.
-In Fig.
-4, tokens highlighted in red in the top 5 most attended by the UT/MT-Encoders relate to molecular weight or polar functional groups, indicating the model does pay attention to key textual information.
-Also, the two molecular graphs plotted using the most attended molecular tokens by the MT-Encoder under different prompts show the original prompt leads to more attention on polar functional groups like ester and ketone, meaning critical information is better captured.
+We analyze the attention matrices from the last layers of the UT/MT-Encoder to investigate whether MolTailor pays attention to key information.
+We select the ESOL dataset, which is related to solubility, to conduct experiments.
+If the model attends to information such as molecular weight or polar functional groups that are critical determinants of solubility, it suggests key information is captured.
+In detail, We pass the SMILES from the test set through MolTailor trained on ESOL to obtain the required attention matrices $\boldsymbol{M}_{UT}$ and $\boldsymbol{M}_{MT}$.
+Then, we take the values of the matrix between the “[CLS]” token and the other tokens as the required attention weights.
+The analysis results in Fig.
+4 show that MolTailor does pay attention to solubility-related information, such as “groups”, “ethers”, and other highlighted tokens.
+Notably, the tokens here are merged to form complete words.
+Additionally, the model also pays attention, as seen in the left molecule attention graph, to polar functional groups within the SMILES.
+Furthermore, variations between the two molecule attention graphs under different prompts indicate effective prompts help the model better focus on key information that assist in solving the task
 Figure 4: Visualization of the attention, answering Q6.
@@
 Conclusion and Future Work
-Overall, this work provides a new perspective on molecular representation learning, not only striving to include more comprehensive information in the representations, but also combining contextual information and obtaining molecular representations more suitable for specific tasks through tailoring.
-In this work, we not only utilize the knowledge implied in the text modality, more importantly, we try to leverage the excellent reasoning ability of language models, which will become more important in the age of large models.
+Overall, in this work we propose a new perspective on molecular representation learning.
+Instead of trying to incorporate more information into the representations, we obtain more task-specific representations by combining contextual information.
+At the same time, we not only utilize the knowledge contained in the text modality, but also try to leverage the reasoning ability of language models, which has huge potential in the era of large language models.
 In the future, we first plan to explore new pretraining tasks that can stably improve model performance on both classification and regression tasks.
@@
 @ on our research.
+Appendix A Detailed Experimental Results
+Due to space constraints in the main text, we omitted the standard deviations of the results from downstream experiments, providing only the average values from three trials.
+Tab.
+5 displays the complete experimental results of our method and baselines on four classification tasks in MoleculeNet.
+Additionally, we employed $\Delta\mathrm{AP}$ to assess model performance, as utilized by Seidl et al.
+(2023).
+$\Delta\mathrm{AP}$ quantifies the overall effectiveness of a model by measuring the discrepancy between its ability to correctly identify classes (as captured by average precision) and the inherent frequency of those classes within the dataset (reflected by the base rate), calculated as follows:
+$\Delta\mathrm{AP}=\frac{1}{N}\sum_{i=1}^{N}(\mathrm{AP}_{i}-\mathrm{BaseRate}_{i})$ (10)
+where, $N$ is the number of classes.
+To compute $\Delta\mathrm{AP}$, we first define the average precision ($\mathrm{AP}$) for class $i$:
+$\mathrm{AP}_{i}=\sum_{k=1}^{M}(\mathrm{P}(k)\cdot\Delta r(k))$ (11)
+[…]
+Furthermore, the base rate for class $i$, $\mathrm{BaseRate}_{i}$, is established by:
+$\mathrm{BaseRate}_{i}=\frac{\sum_{j=1}^{M}y_{ij}\cdot\mathrm{mask}_{ij}}{\sum_{j=1}^{M}\mathrm{mask}_{ij}}$ (12)
+where, $y_{ij}$ represents the binary label for the $j^{th}$ sample under class $i$.
+The term $\mathrm{mask}_{ij}$ indicates whether the value is valid.
+The complete code for calculating $\Delta\mathrm{AP}$ is available in the publicly shared code of this project under the class “DeltaAveragePrecision”.
+Tab.
+6 presents the performance of models on regression tasks.
+Furthermore, we supplemented the performance of two recent molecular-text multimodal methods: TCT5 (Christofidellis et al.
+2023) and CLAMP (Seidl et al.
+2023).
+Appendix B Data Overlap Analysis
+To demonstrate that the performance improvements brought by MolTailor were not caused by data leakage, we analyzed the overlap between the pre-training dataset MT-MTR and the downstream task datasets, as shown in Tab.
+7.
+The results indicate that there indeed was some data overlap.Following this, we delved deeper into our analysis by constructing a pre-training dataset MT-MTR, which excluded the overlapping data.
+Subsequently, we trained MolTailor using the newly formed dataset, and the final experimental results are presented in Tab.
+7.
+The results reveal that there is no clear correlation between the degree of data overlap and the model’s performance.
+Moreover, MolTailor still achieved excellent performance on regression tasks, though its performance on classification tasks was impaired.
+We believe this distinction underscores the dual impact of MT-MTR: it enhances the model’s performance on regression tasks while conversely affecting its performance on classification tasks.
+Appendix C Prompt Examples
+Due to policy adjustments at OpenAI, the prompts used to generate pre-training corpora and downstream task descriptions have become unavailable, as shown in Tab.
+1.
+Therefore, in this section, we only present data examples from MT-MTR and descriptions corresponding to downstream tasks.
+Tab.
+LABEL:tab:mt-mtr-example displays five data examples from MT-MTR, while Tab.
+LABEL:tab:molnet-example presents descriptions of eight tasks from the MoleculeNet dataset.
+Appendix D Statistics of Label Counts in MT-MTR
+When constructing MT-MTR, it is theoretically randomized to select 5-10 properties from the 209 molecular properties provided by RDKit to generate task descriptions.
+Therefore, in theory, the number of the 209 properties in the regression labels of MT-MTR should be approximately equal.
+However, due to the high probability of properties starting with “fr”, indicating the count of certain groups, being zero, we limited the selection probability of properties beginning with “fr” during the selection process, as detailed in the file “03-generate-descriptions.py” on GitHub.
+Additionally, we filtered out properties with excessively high absolute values or those that are NaN.
+Following these filters, the statistical situation of the regression labels included in MT-MTR is shown in Fig.
+5.
+The statistical results reveal that, excluding properties beginning with ’fr’, the occurrence probabilities of various properties are relatively similar, which aligns with expectations.
+Dataset BBBP ClinTox HIV Tox21
+#Molecules 2039 1478 41127 7831
+#Split Scaffold Random Scaffold Random
+#Tasks 1 2 1 12
+#Metrics ROC-AUC $\Delta$AP ROC-AUC $\Delta$AP ROC-AUC $\Delta$AP ROC-AUC $\Delta$AP
+Random 48.38±3.43 0.83±2.51 56.01±4.79 2.09±0.48 49.54±1.27 -0.20±0.10 51.11±1.86 0.98±0.32
+RDKit-DP 78.25±3.44 13.14±2.78 67.36±5.62 8.36±3.00 70.85±2.73 22.96±3.49 65.61±2.00 17.16±1.77
+RDKit-FP 87.65±1.70 20.40±0.95 57.13±2.68 3.00±1.54 78.66±0.88 34.32±3.35 76.14±0.96 32.39±0.89
+MACCS-FP 81.64±1.97 17.22±1.73 83.05±13.13 18.53±9.54 77.53±1.29 23.09±2.15 77.27±0.71 27.19±1.04
+Morgan-FP 82.73±4.46 17.14±2.86 72.61±5.94 12.92±3.04 82.65±0.26 41.09±3.17 75.29±1.49 33.77±0.84
+Grover 79.83±1.48 15.87±1.91 87.75±8.38 20.60±9.74 77.47±2.70 25.42±3.66 79.61±0.37 25.55±3.67
+MolCLR 81.27±4.57 16.70±2.32 78.15±2.29 11.03±1.44 71.48±5.56 13.00±3.73 75.61±1.62 22.18±1.06
+Mole-BERT 82.70±2.72 17.18±2.48 81.82±4.62 15.72±1.65 79.35±1.84 25.40±4.27 84.20±0.71 35.77±0.70
+Uni-Mol 79.52±1.97 16.35±0.83 88.65±2.09 23.75±4.32 74.18±3.05 22.54±5.90 78.08±1.60 26.11±1.48
+BioLinkBERT 83.81±3.54 17.86±1.65 87.75±6.30 21.29±5.09 71.24±5.20 22.36±7.71 73.81±1.20 18.69±1.12
+PubMedBERT 89.10±1.73 20.74±0.28 84.29±6.79 18.91±3.97 72.30±4.91 20.44±6.26 73.77±1.17 19.55±1.11
+ChemBERTa-2 84.70±2.17 18.39±2.53 84.21±0.61 13.41±3.36 78.88±1.19 32.40±6.44 80.75±1.74 30.90±1.40
+CHEM-BERT 84.10±2.25 16.98±2.77 93.80±1.11 26.82±2.56 76.99±3.03 26.96±5.93 80.54±1.04 28.75±0.45
+KCL 76.86±1.28 12.99±1.88 60.80±13.17 4.99±3.55 68.48±3.89 11.04±4.04 74.98±2.08 18.34±2.50
+KV-PLM 86.36±1.38 19.24±1.35 81.20±8.53 15.05±6.80 73.52±3.94 21.52±6.17 74.62±1.22 21.55±3.10
+MoMu-ME 80.41±2.91 16.43±1.95 67.99±8.41 6.44±2.16 71.91±1.95 12.16±2.63 74.75±2.26 19.72±1.08
+MoMu-TE 82.24±2.35 17.57±1.29 81.94±3.01 12.54±2.53 67.88±3.25 13.37±3.15 73.07±1.16 20.32±0.22
+MolT5 86.77±2.68 19.62±1.77 89.65±7.45 25.07±11.90 77.14±4.89 27.55±9.37 80.66±1.28 31.69±1.83
+TCT5 85.58±2.51 18.51±18.51 88.73±5.55 22.77±10.73 79.33±1.22 28.30±4.84 81.72±1.61 32.44±1.80
+CLAMP 83.08±0.80 16.40±0.64 83.02±2.85 15.98±2.00 93.38±0.79 60.75±2.55 85.14±1.67 43.67±2.02
+MolTailor 81.15±2.15 15.94±1.84 92.37±4.45 29.58±2.70 77.42±2.33 26.23±4.87 80.67±1.36 30.57±0.81
+MolTailor 84.65±2.12 18.31±1.01 85.95±6.67 19.88±7.53 76.42±2.63 24.77±4.50 80.32±1.59 31.10±1.72
+MolTailor∗∗ 83.08±1.09 17.43±0.30 83.65±5.23 20.11±5.96 76.01±1.30 26.55±2.59 80.90±1.91 31.38±1.02
+Table 5: Experimental results on classification tasks.
+All experiments were conducted under the linear probe setting, and for each method across each task, the learning rate was searched ten times using Optuna.
+We conducted repeated experiments using three non-specifically set random seeds: 1236, 1237, 1238, and reported the mean and standard deviation of these three trials.
+TCT5 (Christofidellis et al.
+2023) and CLAMP (Seidl et al.
+2023) are additionally included as two new multimodal methods.
+Additionally, we employed $\Delta$AP to assess model performance, as utilized by Seidl et al.
+(2023).
+MolTailor denotes the use of PubMedBERT and CHEM-BERT as the backbone, MolTailor represents the use of PubMedBERT and ChemBERTa as the backbone, and MolTailor∗∗ indicates the use of BioLinkBERT and CHEM-BERT as the backbone.
+Bold results indicate the best outcomes, while underlined results signify that MolTailor shows an improvement over the molecular backbone used.
+Dataset ESOL FreeSolv Lip QM8 Params
+#Molecules 1128 642 4200 21786
+#Split Random Random Random Random
+#Tasks 1 1 2 16
+#Metrics RMSE RMSE RMSE MAE
+Random 3.3358±0.0904 5.4831±0.6031 1.3813±0.0170 0.0320±0.0004 -
+RDKit-DP 4.8940±4.9635 2.8068±0.9990 0.9963±0.0448 0.0202±0.0006 -
+RDKit-FP 1.0830±0.1084 2.0725±0.1017 0.9007±0.0248 0.0181±0.0005 -
+MACCS-FP 1.0833±0.1146 1.9086±0.2425 0.9886±0.0546 0.0196±0.0005 -
+Morgan-FP 1.2413±0.1281 2.1896±0.4250 0.8196±0.0252 0.0200±0.0006 -
+Grover 0.8977±0.0594 1.9041±0.1993 0.8301±0.0414 0.0184±0.0007 107M
+MolCLR 1.3421±0.0578 3.0436±0.7911 1.0448±0.0288 0.0219±0.0007 2M
+Mole-BERT 1.1379±0.0362 2.3626±0.1184 0.8316±0.0320 0.0221±0.0007 2M
+Uni-Mol 1.0509±0.0801 2.6913±0.4001 1.0363±0.0617 0.0219±0.0005 48M
+BioLinkBERT 1.1739±0.0472 3.1350±0.1649 1.0589±0.0506 0.0234±0.0008 108M
+PubMedBERT 1.0715±0.0095 2.5999±0.4324 1.0851±0.0491 0.0232±0.0007 108M
+ChemBERTa-2 0.8103±0.0530 1.8439±0.5930 0.7948±0.0209 0.0191±0.0008 3M
+CHEM-BERT 0.7973±0.0462 2.0214±0.5824 0.8571±0.0381 0.0215±0.0011 51M
+KCL 0.8728±0.0524 2.7615±0.2813 0.9868±0.0368 0.0225±0.0007 1M
+KV-PLM 1.1785±0.1242 2.8840±0.3591 1.1004±0.0242 0.0233±0.0009 109M
+MoMu-ME 1.4135±0.0504 2.3229±0.0986 0.9835±0.0600 0.0222±0.0007 2M
+MoMu-TE 1.2562±0.1029 3.1480±0.4755 1.0885±0.0116 0.0250±0.0006 109M
+MolT5 0.7867±0.0464 2.0978±0.2920 0.8760±0.8760 0.0204±0.0007 110M
+TCT5 0.8194±0.0356 1.8784±0.2288 0.8269±0.0377 0.0189±0.0006 110M
+CLAMP 0.8012±0.0651 2.1808±0.0604 0.7978±0.0499 0.0205±0.0007 56M
+MolTailor 0.7234±0.0495 1.7881±0.3262 0.8107±0.0160 0.0196±0.0005 161M
+MolTailor 0.7128±0.0459 1.7826±0.3342 0.7848±0.0122 0.0185±0.0005 112M
+MolTailor∗∗ 0.7329±0.0331 1.7845±0.4055 0.8094±0.0034 0.0184±0.0006 112M
+Table 6: Experimental results on regression tasks.
+The experimental setup is identical to classification tasks, as referenced in Tab.
+5.
+TCT5 (Christofidellis et al.
+2023) and CLAMP (Seidl et al.
+2023) are additionally included as two new multimodal methods.
+MolTailor denotes the use of PubMedBERT and CHEM-BERT as the backbone, MolTailor represents the use of PubMedBERT and ChemBERTa as the backbone, and MolTailor∗∗ indicates the use of BioLinkBERT and CHEM-BERT as the backbone.
+Bold results indicate the best outcomes, while underlined results signify that MolTailor shows an improvement over the molecular backbone used.
+Dataset Overlapping Ratio CHEM-BERT MolTailor MolTailor
+Classification ROC-AUC BBBP 1975 909 46.03% 84.10±2.25 81.15±2.15 83.04±0.62
+ClinTox 1459 766 52.50% 93.80±1.11 92.37±4.45 91.89±2.05
+HIV 41127 618 1.50% 76.99±3.03 77.42±2.33 76.09±2.94
+Tox21 7831 3937 50.27% 80.54±1.04 80.67±1.36 80.25±2.13
+Average - - - 83.86 82.90 82.82
+Regression RMSE / MAE ESOL 1117 651 58.28% 0.7973±0.0462 0.7234±0.0495 0.7602±0.0619
+FreeSolv 642 407 63.40% 2.0214±0.5824 1.7881±0.3262 1.7097±0.2594
+Lip 4200 629 14.98% 0.8571±0.0381 0.8107±0.0160 0.8095±0.0364
+QM8 21766 423 1.94% 0.0215±0.0011 0.0196±0.0005 0.0197±0.0008
+Average - - - 0.9243 0.8355 0.8248
+Table 7: Data overlap analysis results.
+We examined the overlap of SMILES between MT-MTR and eight tasks from MoleculeNet, with the analysis results presented in the left half of the table.
+In the three columns under “Overlapping Ratio”, the first column indicates the number of molecules contained in the dataset, the second column shows the number of molecules that overlap with those in MT-MTR, and the third column represents the proportion of overlap.
+It is important to note that due to our preprocessing of SMILES which includes normalization and deduplication, the number of molecules in the table differs from the counts originally reported.
+On the right, we display the performance changes of the models after removing data overlap, where CHEM-BERT denotes the performance of the molecular backbone used, MolTailor refers to the model trained on the original MT-MTR dataset, and MolTailor refers to the model trained on MT-MTR, the newly constructed dataset with overlapping data removed.
+Table 8: Five data examples from MT-MTR.
+MT-MTR Examples
+1 SMILES:
+[…]
+Task Descriptions:
+Based on the analysis of the chemical compound, the following properties are important in solving the task:
+1.
+NumSaturatedHeterocycles: This descriptor provides information about the number of saturated heterocycles present in the compound.
+It helps in understanding the presence of specific ring structures that can be relevant for the task.
+2.
+Chi4n: This descriptor measures the 4th-order molecular connectivity index.
+It indicates the electronic and topological properties of the compound.
+It can be useful in identifying the compound’s reactivity and structural complexity.
+3. VSA_EState5: This descriptor represents the potential electrophilicity of the compound. It provides insight into the compound’s ability to act as an electron acceptor in chemical reactions.
+4.
+VSA_EState7: This descriptor is associated with the electrophilic properties of the compound.
+It helps in understanding the compound’s electron-rich regions and its ability to act as an electron donor in reactions.
+5.
+NumAromaticRings: This descriptor indicates the number of aromatic rings present in the compound.
+It helps in identifying the presence of specific ring structures that contribute to the compound’s aromaticity.
+By considering these properties, we can gain valuable insights into the compound’s structure, reactivity, and presence of important functional groups or ring systems.
+These insights can aid in understanding the compound’s behavior and its potential applications in various chemical tasks.
+2 SMILES:
+[…]
+Task Description:
+Based on the given descriptors, the following analysis can be made:
+1. SlogP_VSA2: This descriptor represents the hydrophobicity of the compound. A higher value indicates a higher tendency for the compound to dissolve in lipid-like environments.
+2.
+fr_C_O: This descriptor counts the number of occurrences of a carbon-oxygen bond in the compound.
+It provides information about the presence of carbonyl groups, which are important in various chemical reactions.
+3.
+TPSA: The Topological Polar Surface Area (TPSA) characterizes the polarity of a compound.
+A larger TPSA value suggests a higher degree of intermolecular interactions, which can affect solubility and permeability.
+4.
+Chi3n: This descriptor calculates the third component of the molecular connectivity index, which is related to three-bond or three-connectivity information.
+It contributes to understanding the molecular architecture and reactivity.
+5.
+NumHAcceptors: This descriptor counts the number of hydrogen bond acceptor groups in the compound.
+It represents the compound’s ability to form hydrogen bonds with other molecules, influencing solvation and chemical interactions.
+6. NumSaturatedCarbocycles: This descriptor counts the number of saturated carbocycles present in the compound. It provides information about the compound’s rigidity and structural characteristics.
+7. fr_Al_OH: This descriptor counts the number of hydroxyl groups attached to an alkene or aromatic ring. It is relevant in the context of reactions involving alcohols and aromatic compounds.
+8.
+PEOE_VSA8: This descriptor represents the partial equalization of orbital electronegativity (PEOE) potential energy (VSA) associated with the eighth bin.
+It contributes to understanding the compound’s electronic characteristics and intermolecular interactions.
+9.
+BCUT2D_MRHI: This descriptor calculates the maximum value of the Burden Cluster-based Van der Waals atomic surface area.
+It provides information about the surface properties of the compound, influencing its solubility and reactivity.
+Overall, analyzing these descriptors can provide valuable insights into the hydrophobicity, reactivity, polarity, hydrogen bonding potential, structural characteristics, and electronic properties of a chemical compound.
+3 SMILES:
+[…]
+Task Description:
+Based on the given descriptors, the following analysis results have been observed for the chemical compound:
+1.
+BCUT2D_MWLOW: This descriptor calculates the molecular weight of the compound.
+It indicates the lower end of the molecular weight range.
+This information helps determine the compound’s size and potential for various chemical reactions.
+2.
+BalabanJ: This descriptor calculates the Balaban connectivity index, which measures the compound’s molecular complexity and how connected its atoms are.
+Higher values indicate more complex and interconnected structures.
+3.
+Chi1v: This descriptor calculates the chi connectivity index, which characterizes the topological features of a molecule.
+It provides information about the compound’s branching and connectivity patterns.
+4.
+NumAliphaticCarbocycles: This descriptor counts the number of aliphatic carbocycles present in the compound.
+It helps identify the presence of cyclic structures in the molecule, which can impact its stability and reactivity.
+5.
+SlogP_VSA12: This descriptor quantifies the topological polar surface area (TPSA) contribution to the compound’s logarithm of the octanol/water partition coefficient (SlogP).
+It provides information about the compound’s hydrophobic/hydrophilic nature.
+6.
+PEOE_VSA13: This descriptor represents the partial charges on the atoms of a molecule, specifically for peripheral oxygen atoms.
+It characterizes the compound’s electronic distribution and potential interactions with other molecules.
+By analyzing these properties, we can gain insights into the compound’s molecular weight, complexity, branching patterns, presence of cyclic structures, hydrophobicity/hydrophilicity, and electronic distribution.
+These findings are essential for understanding the compound’s behavior, stability, and potential reactions in various chemical processes.
+4 SMILES:
+[…]
+Task Description:
+Based on the given descriptors, the analysis of the chemical task indicates that several properties of the compound are influential in solving the task.
+The EState_VSA10 value represents the electronic state of the compound, specifically the tendency to donate or accept electrons.
+A higher value suggests a higher ability to interact with other molecules through electron transfer.
+EState_VSA7 measures the electronic state in terms of the compound’s charge distribution.
+A higher value indicates a greater emphasis on electron density in specific regions, which may affect the compound’s reactivity.
+PEOE_VSA13 and PEOE_VSA14 are related to the partial charges of the compound, specifically in terms of polarizability.
+These values help understand the nature of charge distribution and electron density in the molecule.
+EState_VSA11 represents the electronic state, particularly the compound’s ability to act as an electron donor. A higher value implies a stronger electron-donating capability.
+SlogP_VSA8 measures the lipophilicity or solubility of the compound. A higher value suggests a greater affinity for lipid-based solvents or hydrophobic environments.
+Chi0n denotes the topological index, specifically related to electronegativity. It helps understand the tendency of the compound to attract electrons.
+To effectively solve the chemical task, consideration must be given to these descriptors, which provide crucial insights into the compound’s electronic and physical properties.
+By understanding these characteristics, a comprehensive analysis of the compound’s behavior and potential reactions can be achieved.
+5 SMILES:
+[…]
+Task Description:
+Based on the given descriptors, the analysis of the chemical compound is as follows:
+1. FpDensityMorgan2: This descriptor measures the density of functional groups in the compound. It can provide information about the complexity and reactivity of the compound.
+2.
+NumHeteroatoms: This descriptor counts the number of heteroatoms (atoms other than carbon and hydrogen) present in the compound.
+It helps in understanding the compound’s potential for various reactions and interactions.
+3. SMR_VSA9: This descriptor is related to the surface area of the compound and provides information about its size and shape. It can be useful in predicting the compound’s solubility and stability.
+4.
+Chi1v: This descriptor represents the first molecular connectivity index, which measures the topological structure of the compound.
+It can help in understanding the compound’s chemical behavior, reactivity, and properties.
+5.
+BCUT2D_CHGLO: This descriptor represents the Burden modified characteristic topological index, which measures the charge distribution in the compound.
+It can provide insights into the compound’s polarity and electrostatic interactions.
+6. SlogP_VSA3: This descriptor is related to the compound’s hydrophobicity and lipophilicity. It can provide information about its ability to dissolve in lipid-based environments.
+By considering these descriptors, we can gain valuable insights into the structural and chemical properties of the compound, which can help in identifying its potential uses, reactions, and behavior.
+Table 9: Descriptions for the datasets from MoleculeNet.
+Tasks Descriptions
+BBBP Analyzing the properties related to this task, the following descriptors might be helpful:
+1. MolLogP: This descriptor calculates the lipophilicity of the compound, which can influence its ability to penetrate the blood-brain barrier.
+2. NumHDonors and NumHAcceptors: These descriptors count the number of hydrogen bond donors and acceptors, respectively, and can provide insights into the compound’s solubility and permeability.
+3. TPSA: The topological polar surface area can indicate the compound’s polarity, affecting its interaction with the barrier.
+4. fr_halogen: This descriptor identifies the presence of halogen atoms, which might influence the compound’s reactivity and ability to cross the barrier.
+5. RingCount: The number of rings in a compound can affect its shape and size, influencing its ability to penetrate the barrier.
+6. MolWt: Molecular weight can provide information about the size of the compound, which might be a factor in its ability to cross the blood-brain barrier.
+7. FractionCSP3: This descriptor calculates the fraction of sp3 hybridized carbons, which can provide insights into the compound’s three-dimensional structure and flexibility.
+In summary, these descriptors can provide valuable insights into a compound’s size, shape, polarity, and lipophilicity.
+ClinTox The following descriptors could be instrumental in solving this task:
+1. MolWt: Molecular weight can influence the compound’s bioavailability and potential toxicity.
+2. NumHAcceptors and NumHDonors: These descriptors provide information about hydrogen bonding, which can affect how a compound interacts with biological systems.
+3. TPSA: Topological polar surface area can indicate how a compound might interact with membranes and proteins, influencing its toxicity.
+4. fr_halogen: The presence of halogen atoms can affect the compound’s reactivity and potential toxicity.
+5. fr_nitro: This descriptor identifies nitro groups, often associated with toxic effects.
+6. MolLogP: Lipophilicity, measured by this descriptor, can influence how a compound is distributed in the body, affecting its toxicity.
+7. NumAromaticRings: Aromatic rings can play a role in the compound’s stability and reactivity, influencing its toxic effects.
+8. fr_urea, fr_sulfide, fr_ketone: These functional group descriptors can provide insights into specific chemical reactivity that might be associated with toxicity.
+In summary, these descriptors encompass various aspects of a compound’s size, shape, polarity, lipophilicity, and functional groups.
+HIV Analyzing the task, the following descriptors could be particularly helpful in solving the task:
+1. MolWt: Molecular weight can be indicative of the compound’s ability to penetrate the HIV viral envelope.
+2. NumHAcceptors and NumHDonors: These descriptors can provide insights into the hydrogen bonding capabilities, which might influence binding with HIV proteins.
+3. MolLogP: This descriptor represents the compound’s lipophilicity, which can affect its interaction with the viral membrane.
+4. fr_alkyl_halide, fr_amide, fr_nitro: These functional group descriptors may indicate specific reactivity patterns with HIV enzymes.
+5. RingCount: The number of rings in the compound may influence its conformation and ability to bind to HIV targets.
+6. FractionCSP3: This descriptor calculates the fraction of sp3 hybridized carbons, providing insights into the compound’s stereochemistry, which can be crucial for binding specificity.
+In summary, these descriptors encompass various aspects of molecular structure and reactivity
+Tox21 Analyzing the properties related to this task, the following descriptors could be helpful:
+1. NumHAcceptors: This descriptor counts the number of hydrogen bond acceptors, which can indicate potential biological interactions and toxicity.
+2. NumRadicalElectrons: Counting unpaired electrons can provide insights into the compound’s reactivity, which may be linked to toxicity.
+3. MolLogP: This descriptor calculates the octanol-water partition coefficient, reflecting the compound’s hydrophobicity, which can influence its biological activity and potential toxicity.
+4. fr_halogen: Identifying the presence of halogen functional groups can be crucial, as certain halogenated compounds are known to be toxic.
+5. fr_nitro: This descriptor identifies nitro groups, which can be associated with toxic effects in some compounds.
+6. TPSA: The topological polar surface area can provide information about the compound’s ability to cross biological membranes, influencing its toxicity.
+7. RingCount: The number of rings in a compound can affect its structural stability and interactions with biological targets, potentially impacting toxicity.
+In summary, these descriptors can provide valuable insights into the compound’s structure, reactivity, and interactions
+ESOL Analyzing this task, the following descriptors could be helpful in solving it:
+1. MolWt: Molecular weight can influence solubility, as smaller molecules generally dissolve more easily.
+2. TPSA: The topological polar surface area reflects the compound’s polarity, affecting its interaction with polar solvents like water.
+3. NumHDonors and NumHAcceptors: These descriptors count hydrogen bond donors and acceptors, respectively, and can provide insights into the compound’s ability to form hydrogen bonds with water.
+4. MolLogP: This descriptor measures the compound’s lipophilicity, which can indicate its solubility in water versus organic solvents.
+5. FractionCSP3: The fraction of sp3 hybridized carbons can provide information about the compound’s three-dimensional structure and its potential solubility.
+6. fr_ether, fr_aldehyde, fr_ketone: These descriptors identify specific functional groups that may influence solubility.
+In summary, these descriptors can offer valuable insights into a compound’s solubility in water, considering factors like molecular weight, polarity, hydrogen bonding capability, lipophilicity, and specific functional groups.
+FreeSolv Analyzing the properties related to this task, the following descriptors may be helpful:
+1. MolWt: Molecular weight can influence solubility and interaction with water molecules.
+2. TPSA: The topological polar surface area can provide insights into the polarity of the compound, affecting its solvation.
+3. NumHDonors and NumHAcceptors: These descriptors count the number of hydrogen bond donors and acceptors, respectively, which can influence the compound’s interaction with water.
+4. MolLogP: This descriptor calculates the octanol-water partition coefficient, reflecting the compound’s hydrophobicity or hydrophilicity.
+5. FractionCSP3: The fraction of sp3 hybridized carbons can provide information about the compound’s stereochemistry and potential interactions with water.
+6. fr_ether, fr_aldehyde, fr_ketone: These descriptors identify specific functional groups that may influence solvation behavior.
+In summary, these descriptors can provide valuable insights into the compound’s electronic distribution, polarity, hydrogen bonding capability, and specific functional group presence
+Lip Analyzing this task, the following descriptors could be helpful:
+1. MolLogP: This descriptor calculates the logarithm of the compound’s partition coefficient between n-octanol and water, providing direct insights into the compound’s lipophilicity.
+2. SlogP_VSA1 to SlogP_VSA12: These descriptors represent the surface area of a molecule contributing to its solubility in octanol, which can be correlated with lipophilicity.
+3. NumHDonors and NumHAcceptors: These descriptors count the number of hydrogen bond donors and acceptors, respectively, which can influence the compound’s interaction with lipid bilayers.
+4. fr_alkyl_halide, fr_ether, fr_ester: These descriptors identify specific functional groups that may affect the compound’s lipophilic behavior.
+5. FractionCSP3: This descriptor calculates the fraction of carbon atoms that are sp3 hybridized, providing insights into the compound’s stereochemistry, which may influence lipophilicity.
+In summary, these descriptors can provide valuable information about the compound’s lipophilic nature and interactions with lipid environments
+QM8 Several descriptors can be identified as potentially significant for understanding and solving the task:
+1.
+MaxEStateIndex & MinEStateIndex: These descriptors calculate the maximum and minimum E-State indices of atoms in a compound, respectively.
+They can provide insights into the electronic state and distribution of charge within the molecule.
+2. MolWt & ExactMolWt: Representing the molecular weight and exact molecular weight, these descriptors can give information about the size and composition of the molecule.
+3. NumValenceElectrons: This descriptor counts the number of valence electrons in a compound, which can be crucial for understanding its bonding and reactivity.
+4. TPSA: The topological polar surface area can provide insights into the molecule’s solubility and permeability.
+5.
+EState_VSA1 to EState_VSA11: These descriptors calculate the E-State values for various atom types and can provide information about the electronic distribution and interactions within the compound.
+6.
+VSA_EState1 to VSA_EState9: Representing the van der Waals surface area by estimating the electronic energy of atoms in a molecule, these descriptors can provide insights into the electronic distribution and interactions within the compound.
+7. FractionCSP3: This descriptor calculates the fraction of carbon atoms in the compound that are sp3 hybridized, providing insights into the compound’s stereochemistry.
+8. NumAromaticRings & NumAliphaticRings: These descriptors count the number of aromatic and aliphatic rings, respectively, in a compound, indicating the presence of cyclic structures.
+9. MolLogP: This descriptor provides the predicted octanol-water partition coefficient, which can be crucial for understanding the compound’s solubility and bioavailability.
+10. fr_ketone & fr_sulfide: Identifying the presence of ketone and sulfide functional groups, these descriptors can help in determining the compound’s chemical reactivity and potential interactions.
+In summary, the above-listed descriptors can offer valuable insights into the electronic distribution, reactivity, and other properties of the compounds in the task.
+Understanding these properties can be instrumental in solving the task at hand.
+Figure 5: Statistical results of regression labels in MT-MTR.
+We counted the occurrence of each property within the regression labels, as provided by RDKit.
+The horizontal axis of the graph represents the name of the property, while the vertical axis indicates the number of occurrences of that property.